OpenStudy (help!!!!):

Explain/answer: using only 2, 3, 4 ,5,6 how many numbers can be formed that are between 350 and 700?

OpenStudy (dumbcow):

For each place value there are 5 possible digits, except the hundreds place which restricted to 3,4,5,6 Also for "300" numbers the tens place is restricted to 5,6 multiply possible digits for each place value "300's" --> 2*5 = 10 (10 numbers between 350 and 399) "400 thru 700" ---> 3*5*5 = 75 (75 numbers between 400 and 700) Total numbers = 10+75 = 85